# The Tantalizing Case of the Quasar J09505128 – I. Presentation of the Data and Detailed Exploration of the Binary Supermassive Black Hole Scenario

### v3

@@ 5.5 On the Detectability of Doppler Boosting in the Light Curves
The light curves of J0950 exhibit neither substantial nor periodic variability, as discussed in Section 4. In addition to the tests in that section, we phase-folded the light curves using the inferred 33 yr period of J0950, binned the magnitude values by phase, and calculated the median magnitudes and spread within each phase bin. We fit the phase-folded, binned light curves with a flat line corresponding to the mean of each light curve, and found reduced values 5, consistent with no strong periodic structure. Nonetheless, we sought to estimate whether Doppler boosting could produce variability signatures that are either large enough to eventually emerge in the light curves, or too subtle that they remain undetectable.
@@
Considering a SBHB system, in which a BH carrying the BLR beams light as it orbits, the observed specific flux depends on the angle between the beam’s velocity vector and the observer’s line of sight, through the relativistic Doppler factor , where and . For a single source, the observed flux scales as , where is the spectral index (see, for example, dorazio2023review).
@@
5.6 Speculation: Evolution of J0950 in the GW landscape
Given our findings above, it is interesting to speculate about the expected evolution of the GW signal from a system like J0950 in the characteristic strain versus frequency diagram. Figure 13 illustrates the path of such a system as its eccentricity evolves from 0.65 to approximately zero, assuming that GW emission is the dominant mechanism driving orbital evolution. This assumption does not account for the ongoing accretion in the system, which implies continued interaction with the surrounding gas. Gas dynamics could influence the timescale of orbital decay and may sustain a non-zero eccentricity for an extended period. Figure 13 shows three different harmonics of the GW signal. The steps involved in producing the components of the figure are outlined in Appendix C. Here, we predict the overall evolution of the system across the PTA and LISA bands using theoretical sensitivity curves by hazboun2019 and robson2019, respectively. For the purposes of this illustration we adopt representative values of the mass and mass ratio, as indicated in Figure 13.
@@ 5.6 Speculation: Evolution of J0950 in the GW landscape
Finally, we estimate the evolutionary time scales for the GW signal from the system, adopting a characteristic binary mass of 10. Using the binary decay lifetime from peters1964, it will take 55 Myr for this object to merge. The details of this calculation are given in Appendix C. The time the system will spend in the PTA frequency band spans from when the harmonic (initially the strongest at the system’s current eccentricity, ) first enters the band, to the point where the harmonic intersects the PTA sensitivity curve. The system will spend 54 Myr in this band. Furthermore, it will take a time comparable to its overall decay lifetime of 55 Myr for it to evolve from its current eccentricity to the low-frequency end of the LISA band. Once in the LISA band, it will merge within 1 month.
@@ Acknowledgements
In this appendix, we show spectral decomposition models for the spectra of J0950. Each model in Figure 14 includes the quasar continuum, Fe ii lines, broad and narrow He ii 4687 (a weak line in every spectrum of J0950) and H 4863 emission lines, and the [O iii] 4959,5007 doublet (see Figure 3 for an annotated illustration of these features), while Figure 15 shows models of the isolated broad H lines. Also presented in both figures are noisy versions of the models. These model spectra are referenced in Section 3.2, where we quantify sources of random and systematic errors by cross correlating the broad H emission lines in the full and isolated spectral models.
@@ Appendix B MCMC posterior distributions of orbital parameters from radial velocity fits of J0950
In Section 5.3, we performed MCMC fits to two radial velocity curves of J0950: one based on the absolute shifts of the isolated broad H models, and another based on the absolute shifts measured from the isolated broad H lines of the data. The fitted orbital parameters are the period (), the time of conjunction (), the natural log of the velocity semi-amplitude (ln ), and and . Figure 16 shows the corner plot of the posterior distributions from fitting the model-based velocity curve, overlaid with those from fitting the data-based curve.
@@ Appendix C Evolution of a system like J0950 in the PTA and LISA gravitational-wave landscape
The characteristic strain in the detector frame of an eccentric binary at the th harmonic, can be determined using a simplified expression from breivik2020,
Here, is the chirp mass, is luminosity distance, is the detector frame GW frequency, is an eccentricity enhancement factor, and is a function involving a series of Bessel functions of the first kind, that captures the contribution of each harmonic and eccentricity to the gravitational radiation power spectrum (see peters&mathews1963).
@@
The equations above were applied iteratively over a range of eccentricities at a specific . Once the orbit becomes circular, the harmonic dominates. The system was then evolved from the frequency where the eccentricity is approximately zero to the frequency at merger, maintaining a fixed . Recall that for this exercise, we defined “merger” as the separation at ISCO in Section 5.6.
The timescales provided in Section 5.6 for such a system to merge and its GW signal to be detected by PTA experiments and LISA, were computed using the decay lifetime equation given by peters1964,

### v4

@@ V.5 On the Detectability of Doppler Boosting in the Light Curves
The light curves of J0950 exhibit neither substantial nor periodic variability, as discussed in Section IV. In addition to the tests in that section, we phase-folded the light curves using the inferred 33 yr period of J0950, binned the magnitude values by phase, and calculated the median magnitudes and spread within each phase bin. We fit the phase-folded, binned light curves with a flat line corresponding to the mean of each light curve, and found reduced values 5, consistent with no strong periodic structure. Nonetheless, we sought to estimate whether Doppler boosting could produce variability signatures that are either large enough to eventually emerge in the light curves, or too subtle that they remain undetectable.
@@
Considering a SBHB system, in which a BH carrying the BLR beams light as it orbits, the observed specific flux depends on the angle between the beam’s velocity vector and the observer’s line of sight, through the relativistic Doppler factor , where and . For a single source, the observed flux scales as , where is the spectral index (see, for example, D’Orazio and Charisi, 2023).
@@
V.6 Speculation: Evolution of J0950 in the GW landscape
Given our findings above, it is interesting to speculate about the expected evolution of the GW signal from a system like J0950 in the characteristic strain versus frequency diagram. Figure 13 illustrates the path of such a system as its eccentricity evolves from 0.65 to approximately zero, assuming that GW emission is the dominant mechanism driving orbital evolution. This assumption does not account for the ongoing accretion in the system, which implies continued interaction with the surrounding gas. Gas dynamics could influence the timescale of orbital decay and may sustain a non-zero eccentricity for an extended period. Figure 13 shows three different harmonics of the GW signal. The steps involved in producing the components of the figure are outlined in Appendix C. Here, we predict the overall evolution of the system across the PTA and LISA bands using theoretical sensitivity curves by Hazboun et al. (2019) and Robson et al. (2019), respectively. For the purposes of this illustration we adopt representative values of the mass and mass ratio, as indicated in Figure 13.
@@ V.6 Speculation: Evolution of J0950 in the GW landscape
Finally, we estimate the evolutionary time scales for the GW signal from the system, adopting a characteristic binary mass of 10. Using the binary decay lifetime from Peters (1964), it will take 55 Myr for this object to merge. The details of this calculation are given in Appendix C. The time the system will spend in the PTA frequency band spans from when the harmonic (initially the strongest at the system’s current eccentricity, ) first enters the band, to the point where the harmonic intersects the PTA sensitivity curve. The system will spend 54 Myr in this band. Furthermore, it will take a time comparable to its overall decay lifetime of 55 Myr for it to evolve from its current eccentricity to the low-frequency end of the LISA band. Once in the LISA band, it will merge within 1 month.
@@ Acknowledgements
In this appendix, we show spectral decomposition models for the spectra of J0950. Each model in Figure 14 includes the quasar continuum, Fe ii lines, broad and narrow He ii 4687 (a weak line in every spectrum of J0950) and H 4863 emission lines, and the [O iii] 4959,5007 doublet (see Figure 3 for an annotated illustration of these features), while Figure 15 shows models of the isolated broad H lines. Also presented in both figures are noisy versions of the models. These model spectra are referenced in Section III.2, where we quantify sources of random and systematic errors by cross correlating the broad H emission lines in the full and isolated spectral models.
@@ Appendix B MCMC posterior distributions of orbital parameters from radial velocity fits of J0950
In Section V.3, we performed MCMC fits to two radial velocity curves of J0950: one based on the absolute shifts of the isolated broad H models, and another based on the absolute shifts measured from the isolated broad H lines of the data. The fitted orbital parameters are the period (), the time of conjunction (), the natural log of the velocity semi-amplitude (ln ), and and . Figure 16 shows the corner plot of the posterior distributions from fitting the model-based velocity curve, overlaid with those from fitting the data-based curve.
@@ Appendix C Evolution of a system like J0950 in the PTA and LISA gravitational-wave landscape
The characteristic strain in the detector frame of an eccentric binary at the th harmonic, can be determined using a simplified expression from Breivik et al. 2020,
Here, is the chirp mass, is luminosity distance, is the detector frame GW frequency, is an eccentricity enhancement factor, and is a function involving a series of Bessel functions of the first kind, that captures the contribution of each harmonic and eccentricity to the gravitational radiation power spectrum (see Peters and Mathews 1963).
@@
The equations above were applied iteratively over a range of eccentricities at a specific . Once the orbit becomes circular, the harmonic dominates. The system was then evolved from the frequency where the eccentricity is approximately zero to the frequency at merger, maintaining a fixed . Recall that for this exercise, we defined “merger” as the separation at ISCO in Section V.6.
The timescales provided in Section V.6 for such a system to merge and its GW signal to be detected by PTA experiments and LISA, were computed using the decay lifetime equation given by Peters 1964,